### v1

# Representation theory of skew braces

Abstract.
According to Letourmy and Vendramin, a representation of a skew brace is a pair of representations on the same vector space, one for the additive group and the other for the multiplicative group, that satisfies a certain compatibility condition. Following their definition, we shall develop the theory of representations of skew braces. We show that the analogs of Maschke’s theorem and Clifford’s theorem hold for skew braces. We also study irreducible representations in prime characteristic.
1. Introduction
In the study of set-theoretic solutions to the Yang-Baxter equation, Rump [10] introduced an algebraic object called brace and showed that braces correspond to the non-degenerate involutive solutions. As a generalization of this, Guarnieri and Vendramin [4] later defined skew brace and showed that skew braces correspond to all non-degenerate solutions. One can then understand the non-degenerate solutions by studying skew braces. For example, by [6] it is known that whether a solution is multipermutation is related to the right nilpotency of the associated skew brace.
Let us recall the definition of brace and skew brace.
Definition 1.1.
A skew (left) brace is a set $A=(A,\cdot,\circ)$ equipped with two group operations $\cdot$ and $\circ$ such that the so-called brace relation
$a\circ(b\cdot c)=(a\circ b)\cdot a^{-1}\cdot(a\circ c)$
holds for all $a,b,c\in A$. The groups $(A,\cdot)$ and $(A,\circ)$ are referred to as the additive group and multiplicative group of $A$, respectively. A brace is a skew brace with an abelian additive group. It is easy to see that $(A,\cdot)$ and $(A,\circ)$ share the same identity, which we denote by $1$. Moreover, for each $a\in A$, we shall write $a^{-1}$ for its inverse in $(A,\cdot)$, and $\overline{a}$ for its inverse in $(A,\circ)$.
Skew brace is a ring-like structure where the brace relation resembles the distributive law. In fact, brace was introduced as a generalization of radical rings [10]. At the same time, skew brace may be regarded as a generalization of groups by the following example.
Example 1.2.
Let $(A,\cdot)$ be any group. Then:
  1. $\bullet$
    $(A,\cdot,\cdot)$ is a skew brace;
  2. $\bullet$
    $(A,\cdot,\cdot^{\mbox{\tiny op}})$ is also a skew brace, where $\cdot^{\mbox{\tiny op}}$ denotes the opposite operation of $\cdot$ defined by $a\cdot^{\mbox{\tiny op}}b=b\cdot a$ for all $a,b\in A$.
Skew braces that arise in this way are essentially just groups. A skew brace of the form $(A,\cdot,\cdot)$ and $(A,\cdot,\cdot^{\mbox{\tiny op}})$, respectively, is therefore said to be trivial and almost trivial.
It then seems natural to ask to what extent can concepts and results from group theory be generalized to skew braces. For example, identifying groups with almost trivial skew braces, one can view the binary operation
(1.1) $a*b=a^{-1}\cdot(a\circ b)\cdot b^{-1}$
in skew braces as an analog of the commutator in groups. There are natural analogs of center (e.g. socle and annihilator) and nilpotency (e.g. left, right, and strong nilpotency) for skew braces defined in terms of this operation *. Factorizations of skew braces were considered in [5, 11] and there is an analog of Itô’s theorem for skew braces. Isoclinism of skew braces was defined in [8] and it was shown that every skew brace is isoclinic to a so-called stem skew brace. The theory of Schur covers for finite skew braces was developed in [9] and it was shown that the Schur cover of a finite skew brace is unique up to isoclinism. As one can see from these examples, there are many similarities between groups and skew braces. There is no doubt that the integrations of techniques from group theory can help us better understand skew braces.
In relation to Schur covers, the notion of (complex projective) representations of skew braces was also introduced in [9] (whose definition is based on the work of [13]), but not much was said about them. This inspired us to investigate representations of skew braces further. The purpose of this paper is to develop the theory of representations of skew braces. After introducing the basic definitions and some concrete examples, we shall show that the analogs of Maschke’s theorem (§4) and Clifford’s theorem (§6) hold for skew brace representations. We shall also prove an analog of a result about irreducible representations in prime characteristic (§7).
Throughout this paper, the symbol $k$ denotes a field.
Notation.
Given a skew brace $A=(A,\cdot,\circ)$, we shall define
$\begin{cases}\lambda_{a}:A\longrightarrow A;&\lambda_{a}(b)=a^{-1}\cdot(a\circ b)\\ \lambda_{a}^{\mbox{\tiny op}}:A\longrightarrow A;&\lambda^{\mbox{\tiny op}}_{a}(b)=(a\circ b)\cdot a^{-1}\end{cases}$
for each $a\in A$. They are automorphisms of $(A,\cdot)$ by the brace relation. It is also well-known that the maps
$\begin{cases}\lambda:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);&a\mapsto\lambda_{a}\\ \lambda^{\mbox{\tiny op}}:(A,\circ)\longrightarrow\mathrm{Aut}(A,\cdot);&a\mapsto\lambda_{a}^{\mbox{\tiny op}}\end{cases}$
are group homomorphisms. Sometimes $\lambda$ is called the lambda map of $A$. We note that $\lambda^{\mbox{\tiny op}}$ is the lambda map of the skew brace $A^{\mbox{\tiny op}}=(A,\cdot^{\mbox{\tiny op}},\circ)$, which is also called the opposite skew brace of $A$ as defined by [7].
Let us also recall that given a skew brace $A=(A,\cdot,\circ)$, a subset $I$ of $A$ is an ideal if it is a normal subgroup of both $(A,\cdot)$ and $(A,\circ)$, and $\lambda_{a}(I)\subseteq I$ for all $a\in I$. In this case, clearly $I=(I,\cdot,\circ)$ is a skew brace. We also have $a\cdot I=a\circ I$ for all $a\in A$, so one can form a quotient skew brace $A/I$.
2. Representations of skew braces
In this section, we introduce ways to construct skew brace representations from group representations. We define left regular representation for a skew brace naturally from the left regular representations of its additive and multiplicative groups. We shall also define direct sum, tensor product, and equivalence of skew brace representations.
In what follows, let $A=(A,\cdot,\circ)$ be any skew brace.
Definition 2.1.
A representation of $A$ over $k$ is a triple $(V,\beta,\rho)$, where
  1. (1)
    $V$ is a vector space over $k$;
  2. (2)
    $\beta:(A,\cdot)\longrightarrow\mathrm{GL}(V)$ is a representation of $(A,\cdot)$;
  3. (3)
    $\rho:(A,\circ)\longrightarrow\mathrm{GL}(V)$ is a representation of $(A,\circ)$;
such that the relation
(2.1) $\beta(\lambda^{\mbox{\tiny op}}_{a}(b))=\rho(a)\beta(b)\rho(a)^{-1}$
holds for all $a,b\in A$. We say that $(V,\beta,\rho)$ is finite dimensional if $\dim_{k}(V)$ is finite, and similarly infinite dimensional if $\dim_{k}(V)$ is infinite. The degree of $(V,\beta,\rho)$ is defined to be $\dim_{k}(V)$.
Remark 2.2.
Observe that in the outer semidirect product
$(A,\cdot)\rtimes_{\lambda^{\mbox{\tiny op}}}(A,\circ)$
defined by $\lambda^{\mbox{\tiny op}}$, the conjugation action of $(A,\circ)$ on $(A,\cdot)$ is given by
$(1,a)\cdot(b,1)\cdot(1,a)^{-1}=(\lambda_{a}^{\mbox{\tiny op}}(b),1)$
for any $a,b\in A$. The relation (2.1) is meant to represent the above formula, where one applies $\beta$ to the elements of $(A,\cdot)$, and $\rho$ to those of $(A,\circ)$. This interpretation of (2.1) was explained by Professor Vendramin to the second-named author in an email exchange.
We shall omit the phrase “over $k$” for simplicity. All vector spaces, representations, and linear maps to be considered are over $k$ unless specified.
Example 2.3.
Observe that the relation (2.1) is vacuous when $\beta$ is a trivial homomorphism. Thus, given any representation $\rho:(A,\circ)\longrightarrow\mathrm{GL}(V)$ of the multiplicative group, we get a representation $(V,\beta_{0},\rho)$ of the skew brace $A$ by taking $\beta_{0}:(A,\cdot)\longrightarrow\mathrm{GL}(V)$ to be trivial.
But given a representation $\beta:(A,\cdot)\longrightarrow\mathrm{GL}(V)$ of the additive group, in general there need not exist any $\rho:(A,\circ)\longrightarrow\mathrm{GL}(V)$ such that $(V,\beta,\rho)$ is a representation of $A$. For example, take $\dim_{k}(V)$ to be finite and let $\chi_{\beta}$ be the character of $\beta$. For such a $\rho$ to exist, by the relation (2.1) we must have
(2.2) $\chi_{\beta}(\lambda_{a}^{\mbox{\tiny op}}(b))=\chi_{\beta}(b)$
for all $a,b\in A$. As the next example shows, this does not hold in general.
Example 2.4.
Consider the brace $(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+,\circ)$, where
$\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\circ\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}=\begin{pmatrix}a_{1}+(-1)^{a_{2}}b_{1}\\ a_{2}+b_{2}\end{pmatrix}$
for all $\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right),\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)\in\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ (c.f. [1, Theorem]). Here $p$ is an odd prime. Let $\zeta_{p}\in k$ be a primitive $p$th root of unity (assume it exists). Then, for the representation $\beta:(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+)\longrightarrow k^{\times}$ defined by
$\beta\begin{pmatrix}1\\ 0\end{pmatrix}=\zeta_{p},\,\ \beta\begin{pmatrix}0\\ 1\end{pmatrix}=1,$
we have $\chi_{\beta}(\lambda_{a}^{\mbox{\tiny op}}(b))=\zeta_{p}^{(-1)^{a_{2}}b_{1}}$ and $\chi_{\beta}(b)=\zeta_{p}^{b_{1}}$ for $a=\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right),\,b=\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)$. Since they are not equal for $a_{2}=1,\,b_{1}=1$, we see that (2.2) is not satisfied.
Example 2.5.
Given a representation $\beta:(A,\cdot)\longrightarrow\mathrm{GL}(V)$ of the additive group, by taking $\rho_{0}:(A,\circ)\longrightarrow\mathrm{GL}(V)$ to be trivial, we can get a representation $(V,\beta,\rho_{0})$ of the skew brace $A$ if and only if $\beta(\lambda_{a}^{\mbox{\tiny op}}(b))=\beta(b)$ holds for all $a,b\in A$. The latter is vacuous when $A$ is almost trivial.
Let $A*A$ be the subgroup of $(A,\cdot)$ generated by $a*b$ for $a,b\in A$, where * is the operation (1.1). It is known that $A*A$ is an ideal of $A$, and clearly
$a\circ b\equiv a\cdot b\pmod{A*A}$
for all $a,b\in A$. Thus, the quotient skew brace $A/(A*A)$ is trivial. Let
$\pi:A\longrightarrow A/(A*A);\,\ \pi(a)=a\cdot(A*A)=a\circ(A*A)$
denote the natural quotient map. Then, we can construct representations of the skew brace $A$ by lifting representations of the group
$A/(A*A)=(A/(A*A),\cdot)=(A/(A*A),\circ)$
via $\pi$, as follows. We remark that $A*A=1$ when $A$ is a trivial skew brace.
Example 2.6.
Given any group representation $\overline{\alpha}:A/(A*A)\longrightarrow\mathrm{GL}(V)$, by letting $\alpha=\overline{\alpha}\circ\pi$, we get a representation $(V,\alpha,\alpha)$ of the skew brace $A$. Indeed, since $\pi$ is a homomorphism with respect to both $\cdot$ and $\circ$, the map $\alpha$ is a homomorphism on both $(A,\cdot)$ and $(A,\circ)$. Moreover, we have
$\alpha(\lambda_{a}^{\mbox{\tiny op}}(b))$ $=\alpha((a\circ b)\cdot a^{-1})$
$=\alpha(a\cdot b\cdot a^{-1})$
$=\alpha(a)\alpha(b)\alpha(a)^{-1}$
for all $a,b\in A$ since $\ker(\alpha)$ contains $A*A$. This verifies the relation (2.1).
Let $kA$ be the vector space over $k$ that has the set $A$ as a basis. Let
$\begin{cases}\beta_{\mbox{\tiny reg}}:(A,\cdot)\longrightarrow\mathrm{GL}(kA);&\beta_{\mbox{\tiny reg}}(a)(b)=a\cdot b\mbox{ for all }a,b\in A,\\ \rho_{\mbox{\tiny reg}}:(A,\circ)\longrightarrow\mathrm{GL}(kA);&\rho_{\mbox{\tiny reg}}(a)(b)=a\circ b\mbox{ for all }a,b\in A,\end{cases}$
respectively, denote the left regular representations of $(A,\cdot)$ and $(A,\circ)$. As we shall verify below, the triplet $(kA,\beta_{\mbox{\tiny reg}},\rho_{\mbox{\tiny reg}})$ is a representation of $A$, and we shall refer to it as the left regular representation of $A$.
Proposition 2.7.
The triplet $(kA,\beta_{\mbox{\tiny reg}},\rho_{\mbox{\tiny reg}})$ is a representation of $A$.
Proof.
For any $a,b,c\in A$, we have
$\rho_{\mbox{\tiny reg}}(a)\beta_{\mbox{\tiny reg}}(b)\rho_{\mbox{\tiny reg}}(a)^{-1}(c)$ $=\rho_{\mbox{\tiny reg}}(a)\beta_{\mbox{\tiny reg}}(b)\rho_{\mbox{\tiny reg}}(\overline{a})(c)$
$=a\circ(b\cdot(\overline{a}\circ c))$
$=(a\circ b)\cdot a^{-1}\cdot(a\circ\overline{a}\circ c)$
$=\lambda_{a}^{\mbox{\tiny op}}(b)\cdot c$
$=\beta_{\mbox{\tiny reg}}({\lambda_{a}^{\mbox{\tiny op}}(b))}(c).$
This verifies the relation (2.1) and the claim follows. ∎
It is natural to ask whether $(kA,\beta^{\prime}_{\mbox{\tiny reg}},\rho^{\prime}_{\mbox{\tiny reg}})$ is a representation of $A$, where
$\begin{cases}\beta^{\prime}_{\mbox{\tiny reg}}:(A,\cdot)\longrightarrow\mathrm{GL}(kA);&\beta^{\prime}_{\mbox{\tiny reg}}(a)(b)=b\cdot a^{-1}\mbox{ for all }a,b\in A,\\ \rho^{\prime}_{\mbox{\tiny reg}}:(A,\circ)\longrightarrow\mathrm{GL}(kA);&\rho^{\prime}_{\mbox{\tiny reg}}(a)(b)=b\circ\overline{a}\mbox{ for all }a,b\in A,\end{cases}$
respectively, denote the right regular representations of $(A,\cdot)$ and $(A,\circ)$. By definition, the triplet $(kA,\beta^{\prime}_{\mbox{\tiny reg}},\rho^{\prime}_{\mbox{\tiny reg}})$ is a representation of $A$ if and only if
$\rho_{\mbox{\tiny reg}}^{\prime}(a)\beta^{\prime}_{\mbox{\tiny reg}}(b)\rho_{\mbox{\tiny reg}}^{\prime}(a)^{-1}(c)$ $=((c\circ a)\cdot b^{-1})\circ\overline{a}$
$\beta^{\prime}_{\mbox{\tiny reg}}(\lambda_{a}^{\mbox{\tiny op}}(b))(c)$ $=c\cdot(a\circ b^{-1})\cdot a^{-1}$
give the same element, namely
(2.3) $((c\circ a)\cdot b^{-1})\circ\overline{a}=c\cdot(a\circ b^{-1})\cdot a^{-1}$
for all $a,b,c\in A$. As the next example shows, this does not hold in general.
Example 2.8.
Let $(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+,\circ)$ be the brace in Example 2.4. Taking $c=a$ and converting $\cdot$ to $+$, we may write (2.3) as
$(a\circ a)-b=a\circ(-b)\circ a.$
But letting $a=\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right)$ and $b=\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)$, we have
$(a\circ a)-b$ $=\begin{pmatrix}a_{1}+(-1)^{a_{2}}a_{1}-b_{1}\\ b_{2}\end{pmatrix},$
$a\circ(-b)\circ a$ $=\begin{pmatrix}a_{1}+(-1)^{a_{2}}(-b_{1})+(-1)^{a_{2}+b_{2}}a_{1}\\ b_{2}\end{pmatrix}.$
The first components are not equal when $a_{2}=1,\,b_{2}=0$, and $b_{1}\neq 0$.
One can define direct sums and tensor products of skew brace representations the exact same way as for group representations.
Definition 2.9.
Let $(V_{1},\beta_{1},\rho_{1})$ and $(V_{2},\beta_{2},\rho_{2})$ be representations of $A$.
  1. (1)
    Their direct sum is defined to be $(V_{1}\oplus V_{2},\beta_{1}\oplus\beta_{2},\rho_{1}\oplus\rho_{2})$, where
    $(\beta_{1}\oplus\beta_{2})(a)(v_{1},v_{2})$ $=(\beta_{1}(a)(v_{1}),\beta_{2}(a)(v_{2}))$
    $(\rho_{1}\oplus\rho_{2})(a)(v_{1},v_{2})$ $=(\rho_{1}(a)(v_{1}),\rho_{2}(a)(v_{2}))$
    for all $a\in A$ and $v_{1}\in V_{1},\,v_{2}\in V_{2}$.
  2. (2)
    Their tensor product is defined to be $(V_{1}\otimes_{k}V_{2},\beta_{1}\otimes\beta_{2},\rho_{1}\otimes\rho_{2})$, where
    $(\beta_{1}\otimes\beta_{2})(a)(v_{1}\otimes v_{2})$ $=\beta_{1}(a)(v_{1})\otimes\beta_{2}(a)(v_{2})$
    $(\rho_{1}\otimes\rho_{2})(a)(v_{1}\otimes v_{2})$ $=\rho_{1}(a)(v_{1})\otimes\rho_{2}(a)(v_{2})$
    for all $a\in A$ and $v_{1}\in V_{1},\,v_{2}\in V_{2}$.
For both constructions, the relation (2.1) is clearly satisfied.
A representation of $A$ is essentially a vector space equipped with an $(A,\cdot)$-action and an $(A,\circ)$-action that satisfy the compatibility condition (2.1). It is then natural to define “homomorphism” to be a linear map that preserves both of the actions. More specifically:
Definition 2.10.
Let $(V_{1},\beta_{1},\rho_{1})$ and $(V_{2},\beta_{2},\rho_{2})$ be representations of $A$. A homomorphism from $(V_{1},\beta_{1},\rho_{1})$ to $(V_{2},\beta_{2},\rho_{2})$ is a linear map $f:V_{1}\longrightarrow V_{2}$ such that both of the conditions
$f\circ\beta_{1}(a)$ $=\beta_{2}(a)\circ f$
$f\circ\rho_{1}(a)$ $=\rho_{2}(a)\circ f$
hold for all $a\in A$. An isomorphism is a bijective homomorphism. We shall say that $(V_{1},\beta_{1},\rho_{1})$ and $(V_{2},\beta_{2},\rho_{2})$ are equivalent if there is an isomorphism between them.
Suppose that $A=(A,\cdot,\cdot)$ is a trivial skew brace. Even up to equivalence, a representation $(V,\alpha)$ of the group $(A,\cdot)$ may be identified with the representation $(V,\alpha,\alpha)$ of the skew brace $A$. For any $f\in\mathrm{GL}(V)$, clearly
$\alpha_{f}:(A,\cdot)\longrightarrow\mathrm{GL}(V);\,\ \alpha_{f}(a)=f\circ\alpha(a)\circ f^{-1}$
is a representation of $(A,\cdot)$ that is equivalent to $(V,\alpha)$. Note that $(V,\alpha,\alpha_{f})$ is a representation of $A$ if and only if
$\alpha(a)\alpha(b)\alpha(a)^{-1}=\alpha_{f}(a)\alpha(b)\alpha_{f}(a)^{-1}$
for all $a,b\in A$, or equivalently by replacing $a$ with $a^{-1}$, if the commutator
$\alpha_{f}(a)\alpha(a)^{-1}=f\alpha(a)f^{-1}\alpha(a)^{-1}$
centralizes $\mathrm{Im}(\alpha)$ for all $a\in A$. But $(V,\alpha,\alpha)$ and $(V,\alpha,\alpha_{f})$ are equivalent if and only if $\alpha_{f}=\alpha$. Thus, if we can find $f\in\mathrm{GL}(V)$ such that
  1. (1)
    $f\alpha(a)f^{-1}\alpha(a)^{-1}$ centralizes $\mathrm{Im}(\alpha)$ for all $a\in A$;
  2. (2)
    $f$ does not centralize $\mathrm{Im}(\alpha)$;
then even though $(V,\alpha)$ and $(V,\alpha_{f})$ are equivalent as representations of the group $(A,\cdot)$, the triplets $(V,\alpha,\alpha)$ and $(V,\alpha,\alpha_{f})$ are not equivalent as representations of the skew brace $A$. We shall give a concrete example below. It follows that the equivalence of representations of skew braces is much more complicated than that of representations of groups.
Example 2.11.
Consider the symmetric group $S_{3}=(S_{3},\cdot)$ on three letters. Let $\alpha:S_{3}\longrightarrow\mathrm{GL}_{2}(k)$ be the representation defined by
[…]
Assume $\mathrm{char}(k)\neq 2,3$, and take $f\in\mathrm{GL}_{2}(k)$ to be the matrix
$f=\begin{pmatrix}-1&2\\ -2&1\end{pmatrix},$
which is invertible since $\mathrm{char}(k)\neq 3$. Then $f$ satisfies (1) above because
[…]
3. Irreducible and indecomposable representations
In this section, we define invariant subspaces, irreducibility, and indecomposability for skew brace representations. We give concrete examples to show that irreducible and indecomposable representations of skew braces are way more complicated than those of groups. We shall also prove some simple but fundamental facts about these representations, such as the analog of Schur’s lemma.
In what follows, let $A=(A,\cdot,\circ)$ be any skew brace.
Definition 3.1.
Let $(V,\beta,\rho)$ be a representation of $A$. A subspace $U$ of $V$ is said to be $(\beta,\rho)$-invariant if it is both $\beta$-invariant and $\rho$-invariant. In this case, the triplet $(U,\beta_{U},\rho_{U})$ is also a representation of $A$, where
$\begin{cases}\beta_{U}:(A,\cdot)\longrightarrow\mathrm{GL}(U);&\beta_{U}(a)=\beta(a)|_{U},\\ \rho_{U}:(A,\circ)\longrightarrow\mathrm{GL}(U);&\rho_{U}(a)=\rho(a)|_{U},\end{cases}$
respectively, denote the sub-representations of $\beta$ and $\rho$ restricted to $U$. We shall say that the representation $(V,\beta,\rho)$ is
  1. (1)
    irreducible if $V\neq\{0\}$ and $V$ does not have any non-zero proper $(\beta,\rho)$-invariant subspace;
  2. (2)
    indecomposable if $V\neq\{0\}$ and $V$ cannot be expressed as a direct sum of two non-zero proper $(\beta,\rho)$-invariant subspaces;
  3. (3)
    completely reducible if $V$ can be expressed as a direct sum of irreducible $(\beta,\rho)$-invariant subspaces, namely $(\beta,\rho)$-invariant subspaces $U$ for which the representation $(U,\beta_{U},\rho_{U})$ is irreducible.
The next proposition gives a way to construct a $(\beta,\rho)$-invariant subspace from a $\beta$-invariant or $\rho$-invariant subspace.
Proposition 3.2.
Let $(V,\beta,\rho)$ be a representation of $A$.
  1. $(a)$
    If $U$ is a $\beta$-invariant subspace of $V$, then so is $\rho(a)(U)$ for any $a\in A$, and in particular, we get that the sum
    (3.1) $\sum_{a\in A}\rho(a)(U)$
    is a $(\beta,\rho)$-invariant subspace.
  2. $(b)$
    If $U$ is a $\rho$-invariant subspace of $V$, then the sum
    (3.2) $\sum_{b\in A}\beta(b)(U)$
    is a $(\beta,\rho)$-invariant subspace.
Proof of $(a)$.
Fix $a\in A$. For any $b\in A$ and $u\in U$, the relation (2.1) yields
$\beta(\lambda_{a}^{\mbox{\tiny op}}(b))(\rho(a)(u))=\rho(a)(\beta(b)(u)),$
which lies in $\rho(a)(U)$ because $U$ is $\beta$-invariant. Since $\lambda_{a}^{\mbox{\tiny op}}$ is a bijection on $A$, this shows that $\rho(a)(U)$ is $\beta$-invariant. Thus, the sum (3.1), which is clearly $\rho$-invariant, is also $\beta$-invariant. ∎
Proof of $(b)$.
For any $a,b\in A$ and $u\in U$, the relation (2.1) yields
$\rho(a)(\beta(b)(u))=\beta(\lambda_{a}^{\mbox{\tiny op}}(b))(\rho(a)(u)),$
which lies in $\beta(\lambda_{a}^{\mbox{\tiny op}}(b))(U)$ and hence in (3.2) because $U$ is $\rho$-invariant. Thus, the sum (3.2), which is clearly $\beta$-invariant, is also $\rho$-invariant. ∎
Corollary 3.3.
Let $(V,\beta,\rho)$ be a representation of $A$. The sums
$\sum_{a,b\in A}\rho(a)\beta(b)(U)\mbox{ and }\sum_{a,b\in A}\beta(b)\rho(a)(U)$
are $(\beta,\rho)$-invariant subspaces for any subspace $U$ of $V$.
Proof.
This follows immediately from Proposition 3.2 because
$\sum_{b\in A}\beta(b)(U)\mbox{ and }\sum_{a\in A}\rho(a)(U)$
are clearly $\beta$-invariant and $\rho$-invariant, respectively. ∎
It is well-known that the degree of an irreducible representation of a finite group $G$ is at most $|G|$. A similar statement holds for a finite skew brace.
Proposition 3.4.
Suppose that $A$ is a finite skew brace. Then the degree of an irreducible representation of $A$ is at most $|A|^{2}$.
Proof.
Let $(V,\beta,\rho)$ be an irreducible representation of $A$. Let $v\in V$ be any non-zero vector and let $U=\mathrm{span}_{k}(v)$. The non-zero subspace
$\sum_{a,b\in A}\rho(a)\beta(b)(U),$
which is $(\beta,\rho)$-invariant by Corollary 3.3, equals $V$ because $(V,\beta,\rho)$ is irreducible. Since the above subspace is generated by
$\{\rho(a)\beta(b)(v):a,b\in A\},$
it follows that $\dim_{k}(V)$ is at most $|A|^{2}$. ∎
A representation $(V,\beta,\rho)$ of $A$ is clearly irreducible if either $(V,\beta)$ or $(V,\rho)$ is irreducible as a group representation, and similarly for indecomposable. It is possible, as Examples 3.5 and 3.6 below show, respectively, that
  1. (1)
    both $(V,\beta)$ and $(V,\rho)$ are reducible, yet $(V,\beta,\rho)$ is irreducible;
  2. (2)
    both $(V,\beta)$ and $(V,\rho)$ are decomposable, yet $(V,\beta,\rho)$ is indecomposable though reducible.
Irreducible and indecomposable representations of skew braces are therefore much more difficult to understand than those of groups.
Example 3.5.
Consider the brace $(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+,\circ)$, where
$\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\circ\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}=\begin{pmatrix}a_{1}+b_{1}+a_{2}b_{2}\\ a_{2}+b_{2}\end{pmatrix}$
for all $\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right),\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)\in\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ (c.f. [3, Proposition 2.4]), and note that
$(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},\circ)=\left\langle\begin{pmatrix}1\\ 1\end{pmatrix}\right\rangle=\left\{\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1\\ 1\end{pmatrix},\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\end{pmatrix}\right\}$
is cyclic. Let $\beta:(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+)\longrightarrow\mathrm{GL}_{2}(k)$ be defined by
$\beta\left(\begin{matrix}1\\ 0\end{matrix}\right)=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix},\,\beta\left(\begin{matrix}0\\ 1\end{matrix}\right)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$
and let $\rho:(\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},\circ)\longrightarrow\mathrm{GL}_{2}(k)$ be defined by
$\rho\left(\begin{matrix}1\\ 1\end{matrix}\right)=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}.$
To verify the relation (2.1), note that it may be written as
$\beta\begin{pmatrix}a_{2}b_{2}\\ 0\end{pmatrix}=\rho\left(\begin{matrix}a_{1}\\ a_{2}\end{matrix}\right)\beta\left(\begin{matrix}b_{1}\\ b_{2}\end{matrix}\right)\rho\left(\begin{matrix}a_{1}\\ a_{2}\end{matrix}\right)^{-1}\beta\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}^{-1}.$
For $a_{2}=0$ or $b_{2}=0$, this is trivial since $\beta\left(\begin{smallmatrix}*\\ 0\end{smallmatrix}\right),\rho\left(\begin{smallmatrix}*\\ 0\end{smallmatrix}\right)=\pm\left(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right)$. For $a_{2}=1$ and $b_{2}=1$, this holds because $\beta\left(\begin{smallmatrix}*\\ 1\end{smallmatrix}\right)=\pm\left(\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right),\,\rho\left(\begin{smallmatrix}*\\ 1\end{smallmatrix}\right)=\pm\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$ and
$\beta\begin{pmatrix}1\\ 0\end{pmatrix}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}^{-1}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}^{-1}.$
Thus $(k^{2},\beta,\rho)$ is a representation of the brace under consideration.
Assume that $\mathrm{char}(k)\neq 2$ and that $-1$ has a square root $i$ in $k$. Then
  1. $\bullet$
    the one-dimensional $\beta$-invariant subspaces are $\mathrm{span}_{k}\left(\begin{smallmatrix}1\\ 1\end{smallmatrix}\right),\,\mathrm{span}_{k}\left(\begin{smallmatrix}-1\\ 1\end{smallmatrix}\right)$;
  2. $\bullet$
    the one-dimensional $\rho$-invariant subspaces are $\mathrm{span}_{k}\left(\begin{smallmatrix}i\\ 1\end{smallmatrix}\right),\,\mathrm{span}_{k}\left(\begin{smallmatrix}-i\\ 1\end{smallmatrix}\right)$.
Thus, the group representations $(k^{2},\beta)$ and $(k^{2},\rho)$ are both reducible. Since the four one-dimensional subspaces above are pairwise distinct, there are no one-dimensional $(\beta,\rho)$-invariant subspaces, whence the brace representation $(k^{2},\beta,\rho)$ is irreducible.
Example 3.6.
Let $(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+,\circ)$ be the brace in Example 2.4. Note that $\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)\circ\left(\begin{smallmatrix}x\\ 0\end{smallmatrix}\right)\circ\overline{\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)}=-\left(\begin{smallmatrix}x\\ 0\end{smallmatrix}\right)$ for all $x\in\mathbb{Z}/p\mathbb{Z}$, and we have
$(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},\circ)$ =$=\left\{\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1\\ 0\end{pmatrix},\dots,\begin{pmatrix}p-1\\ 0\end{pmatrix}\right\}\rtimes\left\{\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\end{pmatrix}\right\}.$
Assume $\mathrm{char}(k)=p$. Let $\beta:(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},+)\longrightarrow\mathrm{GL}_{3}(k)$ be defined by
$\beta\begin{pmatrix}1\\ 0\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&1\end{pmatrix},\,\ \beta\begin{pmatrix}0\\ 1\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},$
and let $\rho:(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z},\circ)\longrightarrow\mathrm{GL}_{3}(k)$ be defined by
$\rho\begin{pmatrix}1\\ 0\end{pmatrix}=\begin{pmatrix}1&0&0\\ 1&1&0\\ 0&0&1\end{pmatrix},\,\ \rho\begin{pmatrix}0\\ 1\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}.$
To verify the relation (2.1), note that it may be written as
$\beta\begin{pmatrix}(-1)^{a_{2}}b_{1}\\ b_{2}\end{pmatrix}=\rho\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\beta\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}\rho\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}^{-1}.$
We may put $b_{2}=0$ because $\beta\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ is the identity matrix. Since
$\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}=\begin{pmatrix}0\\ a_{2}\end{pmatrix}\circ\begin{pmatrix}(-1)^{a_{2}}a_{1}\\ 0\end{pmatrix}$
and $\rho\left(\begin{smallmatrix}*\\ 0\end{smallmatrix}\right)$ commutes with elements of $\mathrm{Im}(\beta)$, the value of $a_{1}$ does not matter so we may put $a_{1}=0$. We also see that the relation is trivial for $a_{2}=0$. For $b_{2}=0,\,a_{1}=0$, and $a_{2}=1$, in matrix form the relation becomes
$\begin{pmatrix}1&0&0\\ 0&1&-b_{1}\\ 0&0&1\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&b_{1}\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix}^{-1},$
which is easily verified to be true. Hence, the triplet $(k^{3},\beta,\rho)$ is a representation of the brace under consideration.
To simplify notation, let $e_{1},\,e_{2},\,e_{3}$ denote the standard basis vectors of $k^{3}$. The group representation $(k^{3},\beta)$ is decomposable because
$k^{3}=\mathrm{span}_{k}(e_{1})\oplus\mathrm{span}_{k}(e_{2},e_{3})$
and the two direct summands are $\beta$-invariant. Similarly, the group representation $(k^{3},\rho)$ is also decomposable because
$k^{3}=\mathrm{span}_{k}(e_{1},e_{2})\oplus\mathrm{span}_{k}(e_{3})$
and the two direct summands are $\rho$-invariant. A computation of eigenspaces yields that $\mathrm{span}_{k}(e_{2})$ is the only one-dimensional $(\beta,\rho)$-invariant subspace of $k^{3}$. Thus, if $(k^{3},\beta,\rho)$ were decomposable, then necessarily
$k^{3}=\mathrm{span}_{k}(e_{2})\oplus U$
for some $(\beta,\rho)$-invariant subspace $U$. But for any $u\in U$, we have
$\begin{cases}\beta\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)(u)-u=u_{3}e_{2}\\[4.0pt] \rho\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)(u)-u=u_{1}e_{2}\end{cases}\mbox{ for }u=\begin{pmatrix}u_{1}\\ u_{2}\\ u_{3}\end{pmatrix}.$
Since $\mathrm{span}_{k}(e_{2})\cap U=\{0\}$, we get $u_{1}=u_{3}=0$ and $U$ would be contained in $\mathrm{span}_{k}(e_{2})$, which is impossible. This shows that $(V,\beta,\rho)$ is indecomposable, though it is reducible since $\mathrm{span}_{k}(e_{2})$ is a $(\beta,\rho)$-invariant subspace.
Let us also give an example of a representation $(V,\beta,\rho)$ of $A$ where $(V,\beta)$ is irreducible, while $(V,\rho)$ is reducible but indecomposable.
Example 3.7.
Consider the skew brace $(S_{3},\cdot,\circ)$, where
$\sigma\circ\tau=\sigma_{1}\tau\sigma_{2}$
[…]
Assume $\mathrm{char}(k)=2$. Let $\beta:(S_{3},\cdot)\longrightarrow\mathrm{GL}_{2}(k)$ be defined by
[…]
and let $\rho:(S_{3},\circ)\longrightarrow\mathrm{GL}_{2}(k)$ be defined by
[…]
To verify the relation (2.1), note that it may be written as
$\beta(\sigma_{1})\beta(\tau)\beta(\sigma_{1})^{-1}=\rho(\sigma)\beta(\tau)\rho(\sigma)^{-1},$
[…]
It is clear that the group representation $(k^{2},\beta)$ is irreducible, while $(k^{2},\rho)$ is reducible but indecomposable. The skew brace representation $(k^{2},\beta,\rho)$ is of course also irreducible.
Next, we show that the analog of Schur’s lemma holds for skew brace representations. The proof is essentially the same as the case of groups, but we still include it for the sake of completeness.
Lemma 3.8.
Let $(V_{1},\beta_{1},\rho_{1}),\,(V_{2},\beta_{2},\rho_{2})$ be irreducible representations of $A$ and let $f:V_{1}\longrightarrow V_{2}$ be a homomorphism. Then $f$ is either the zero map or an isomorphism.
Proof.
It is clear that $\ker(f)$ is a $(\beta_{1},\rho_{1})$-invariant subspace of $V_{1}$ and $\mathrm{Im}(f)$ is a $(\beta_{2},\rho_{2})$-invariant subspace of $V_{2}$. Thus, if $f$ is not the zero map, then by irreducibility, we have $\ker(f)=\{0\}$ and $\mathrm{Im}(f)=V_{2}$, which means that $f$ is an isomorphism. ∎
Proposition 3.9.
Let $(V,\beta,\rho)$ be a finite dimensional irreducible representation of $A$ and let $f:V\longrightarrow V$ be an endomorphism. If $k$ is algebraically closed, then $f$ is given by scalar multiplication.
Proof.
Let $\lambda$ be any eigenvalue of $f$. Clearly $f-\lambda\cdot\mathrm{id}_{V}$ has non-trivial kernel and is also a homomorphism. Thus, by Lemma 3.8, it must be the zero map, namely $f$ is scalar multiplication by $\lambda$. ∎
There is a similar result regarding indecomposable representations and its analog is also true for skew brace representations. Again, the proof is essentially the same as the case of groups.
Recall that a linear transformation $f:V\longrightarrow V$ on a vector space $V$ is said to be nilpotent if $f^{n}=0_{V}$ for some natural number $n$.
Proposition 3.10.
Let $(V,\beta,\rho)$ be a finite dimensional indecomposable representation of $A$ and let $f:V\longrightarrow V$ be an endomorphism. If $k$ is algebraically closed, then $f$ is either nilpotent or an isomorphism.
Proof.
Since $V$ is a direct sum of the generalized eigenspaces of $f$, all of which are plainly $(\beta,\rho)$-invariant, there can only be one eigenvalue $\lambda$ of $f$. Then $f$ is an isomorphism when $\lambda\neq 0$, and is nilpotent otherwise. ∎
4. Analog of Maschke’s theorem
Let us first recall Maschke’s theorem for group representations. We recall the proof as well because we shall need the idea involved.
Theorem 4.1 (Maschke’s theorem).
Let $G$ be a finite group and let $k$ be a field whose characteristic is 00 or coprime to $|G|$. For any finite dimensional representation $(V,\beta)$ of $G$ over $k$, every $\beta$-invariant subspace of $V$ admits a $\beta$-invariant complement in $V$. Therefore, by induction on the degree, every finite dimensional representation of $G$ over $k$ is completely reducible.
Proof.
Let $U$ be any $\beta$-invariant subspace of $V$. Then $V=U\oplus U^{\prime}$ for some subspace $U^{\prime}$ because $\dim_{k}(V)$ is finite. Here $U^{\prime}$ need not be $\beta$-invariant. Let
$\pi:V\longrightarrow U;\,\ \pi(u+u^{\prime})=u\,\ (u\in U,\,u^{\prime}\in U^{\prime})$
denote the natural projection map, and consider the linear map
$f:V\longrightarrow V;\,\ f=\frac{1}{|G|}\sum_{g\in G}\beta(g)\circ\pi\circ\beta(g)^{-1},$
where $|G|$ is invertible in $k$ by the hypothesis. It is straightforward to check that $f\circ f=f$ and $\mathrm{Im}(f)=U$. Thus, we have
$V=\mathrm{Im}(f)\oplus\ker(f)=U\oplus\ker(f).$
It is easy to verify that $f$ preserves the action of $G$. This implies that $\ker(f)$ is a $\beta$-invariant subspace, whence the claim. ∎
The analog of Maschke’s theorem for skew brace representations holds.
Theorem 4.2 (Analog of Maschke’s theorem).
Let $A=(A,\cdot,\circ)$ be a finite skew brace and let $k$ be a field whose characteristic is 00 or coprime to $|A|$. For any finite dimensional representation $(V,\beta,\rho)$ of $A$ over $k$, every $(\beta,\rho)$-invariant subspace of $V$ admits a $(\beta,\rho)$-invariant complement in $V$. Therefore, by induction on the degree, every finite dimensional representation of $A$ over $k$ is completely reducible.
Proof.
Let $U$ be any $(\beta,\rho)$-invariant subspace of $V$. From Theorem 4.1, we know that $V=U\oplus U^{\prime}$ for some $\rho$-invariant subspace $U^{\prime}$. Here $U^{\prime}$ need not be $\beta$-invariant. Let $\pi:V\longrightarrow U$ be the natural projection map and let
$f:V\longrightarrow V;\,\ f=\frac{1}{|A|}\sum_{a\in A}\beta(a)\circ\pi\circ\beta(a)^{-1}$
be the linear map as in the proof of Theorem 4.1. We know that $\ker(f)$ is a $\beta$-invariant subspace such that $V=U\oplus\ker(f)$. Below, we shall show that $\rho(a)\circ f=f\circ\rho(a)$ for all $a\in A$, so that $\ker(f)$ is also $\rho$-invariant. The claim then follows because $\ker(f)$ is a $(\beta,\rho)$-invariant complement of $U$ in $V$.
Let $a\in A$. First we use the relation (2.1) to compute that
$\rho(a)\circ f$ $=\frac{1}{|A|}\sum_{b\in A}\rho(a)\circ\beta(b)\circ\pi\circ\beta(b)^{-1}$
$=\frac{1}{|A|}\sum_{b\in A}\beta(\lambda_{a}^{\mbox{\tiny op}}(b))\circ\rho(a)\circ\pi\circ\beta(b)^{-1}.$
Note that $\rho(a)$ and $\pi$ commute since $U$ and $U^{\prime}$ are both $\rho$-invariant. Again, using the relation (2.1) but with $b$ replaced by $b^{-1}$, we obtain
$\rho(a)\circ f$ $=\frac{1}{|A|}\sum_{b\in A}\beta(\lambda_{a}^{\mbox{\tiny op}}(b))\circ\pi\circ\rho(a)\circ\beta(b)^{-1}$
$=\frac{1}{|A|}\sum_{b\in A}\beta(\lambda_{a}^{\mbox{\tiny op}}(b))\circ\pi\circ\beta(\lambda_{a}^{\mbox{\tiny op}}(b))^{-1}\circ\rho(a)$
$=\left(\frac{1}{|A|}\sum_{b\in A}\beta(\lambda_{a}^{\mbox{\tiny op}}(b))\circ\pi\circ\beta(\lambda_{a}^{\mbox{\tiny op}}(b))^{-1}\right)\circ\rho(a)$
$=f\circ\rho(a),$
where the last equality holds because $\lambda_{a}^{\mbox{\tiny op}}$ is a bijection on $A$. ∎
It is well-known that Maschke’s theorem admits a converse, as follows. We give the proof since we shall need the idea involved.
Theorem 4.3 (Converse of Maschke’s theorem).
Let $G$ be a finite group and let $k$ be a field such that for any finite dimensional representation $(V,\beta)$ of $G$ over $k$, every $\beta$-invariant subspace of $V$ admits a $\beta$-invariant complement in $V$. Then the characteristic of $k$ is 00 or coprime to $|G|$.
Proof.
Consider the subspace $\mathrm{span}_{k}(\delta)$ generated by $\delta=\sum_{g\in G}g$ in the left regular representation $(kG,\beta_{\mbox{\tiny reg}})$ of $G$. Clearly it is $\beta_{\mbox{\tiny reg}}$-invariant, so
$kG=\mathrm{span}_{k}(\delta)\oplus U$
for some $\beta_{\mbox{\tiny reg}}$-invariant subspace $U$. But $U$ must be equal to the subspace
$U^{\prime}=\Bigg{\{}\sum_{g\in G}c_{g}g\in kG:\sum_{g\in G}c_{g}=0\Bigg{\}}.$
Indeed, for any $\sum_{g\in G}c_{g}g\in U$, the $\beta_{{\mbox{\tiny reg}}}$-invariance of $U$ implies that
$\sum_{h\in G}\beta_{\mbox{\tiny reg}}(h)\Bigg{(}\sum_{g\in G}c_{g}g\Bigg{)}=\sum_{g\in G}c_{g}\sum_{h\in G}hg=\Bigg{(}\sum_{g\in G}c_{g}\Bigg{)}\delta$
also lies in $U$. Since $\mathrm{span}_{k}(\delta)\cap U=\{0\}$, we must have $\sum_{g\in G}c_{g}=0$, which means that $U\subseteq U^{\prime}$. Since $U$ and $U^{\prime}$ both have dimension $|G|-1$, we have equality $U=U^{\prime}$. It follows that $\mathrm{char}(k)$ is 00 or coprime to $|G|$, for otherwise $\delta\in U^{\prime}$ and $U^{\prime}$ cannot be a complement of $\mathrm{span}_{k}(\delta)$. ∎
The analog of the converse of Maschke’s theorem for skew brace representations is also true. This means that in Theorem 4.2, the assumption on the characteristic of $k$ cannot be dropped for any finite skew brace $A$.
Theorem 4.4 (Analog of converse of Maschke’s theorem).
Let $A=(A,\cdot,\circ)$ be a finite skew brace and let $k$ be a field such that for any finite dimensional representation $(V,\beta,\rho)$ of $A$ over $k$, every $(\beta,\rho)$-invariant subspace of $V$ admits a $(\beta,\rho)$-invariant complement in $V$. Then the characteristic of $k$ is 00 or coprime to $|A|$.
Proof.
Consider the subspace $\mathrm{span}_{k}(\delta)$ generated by $\delta=\sum_{a\in A}a$ in the left regular representation $(kA,\beta_{\mbox{\tiny reg}},\rho_{\mbox{\tiny reg}})$ of $A$. Clearly it is $(\beta_{\mbox{\tiny reg}},\rho_{\mbox{\tiny reg}})$-invariant, so
$kA=\mathrm{span}_{k}(\delta)\oplus U$
for some $(\beta_{\mbox{\tiny reg}},\rho_{\mbox{\tiny reg}})$-invariant subspace $U$. But then $U$ is in particular a $\beta_{\mbox{\tiny reg}}$-invariant subspace, so from the proof of Theorem 4.3, we see that $\mathrm{char}(k)$ is not a prime divisor of $|A|$. ∎
We end with an example that shows that when $\mathrm{char}(k)$ divides $|A|$, a finite dimensional representation of $A$ over $k$ need not be completely reducible.
Example 4.5.
Consider the brace $A=(\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z},+,\circ)$, where
$\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\circ\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}=\begin{pmatrix}a_{1}+b_{1}+a_{2}b_{2}\\ a_{2}+b_{2}\end{pmatrix}$
for all $\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right),\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)\in\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ (c.f. [3, Proposition 2.4]). Here $p$ is a prime. Clearly $A*A=\mathbb{Z}/p\mathbb{Z}\times\{0\}$. Assume $\mathrm{char}(k)=p$. Let $\alpha:A\longrightarrow\mathrm{GL}_{2}(k)$ be the lift of the group representation $\overline{\alpha}:A/A*A\longrightarrow\mathrm{GL}_{2}(k)$ defined by
$\overline{\alpha}\left(\begin{pmatrix}0\\ 1\end{pmatrix}+(A*A)\right)=\overline{\alpha}\left(\begin{pmatrix}0\\ 1\end{pmatrix}\circ(A*A)\right)=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}.$
Then $(k^{2},\alpha,\alpha)$ is a representation of the brace $A$ by Example 2.6, and it is clearly not completely reducible because $\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$ is not diagonalizable.
5. Representations as modules
In this section, as in the case of groups, we reinterpret skew brace representations in the language of modules. This has the advantage that one can employ tools from module theory. For example, we shall prove an analog of the Krull-Schmidt theorem for skew brace representations.
In what follows, let $A=(A,\cdot,\circ)$ be any skew brace.
Definition 5.1.
A (left) $kA$-module is a vector space $V$ over $k$ such that
  1. (1)
    $V$ is a left $k(A,\cdot)$-module, where we write $\cdot$ for the action of $(A,\cdot)$;
  2. (2)
    $V$ is a left $k(A,\circ)$-module, where we write $\circ$ for the action of $(A,\circ)$;
and the two actions satisfy the relation
(5.1) $\lambda_{a}^{\mbox{\tiny op}}(b)\cdot v=a\circ(b\cdot(\overline{a}\circ v))$
for all $a,b\in A$ and $v\in V$.
Clearly there is a one-to-one correspondence between representations of $A$ and $kA$-modules. Precisely, a representation $(V,\beta,\rho)$ of $A$ may be made into a $kA$-module by defining the actions to be
(5.2) $\begin{cases}a\cdot v=\beta(a)(v)\\ a\circ v=\rho(a)(v)\end{cases}$
for all $a\in A$ and $v\in V$, and vice versa. Note that (5.1) is simply a restatement of the relation (2.1) under this correspondence.
The following are counterparts of Examples 2.3 and 2.5.
Example 5.2.
Given any $k(A,\circ)$-module, we can turn it into a $kA$-module by letting $(A,\cdot)$ act trivially.
Example 5.3.
Given a $k(A,\cdot)$-module, by letting $(A,\circ)$ act trivially, we can make it into a $kA$-module if and only if the $(A,\cdot)$-actions of $\lambda_{a}^{\mbox{\tiny op}}(b)$ and $b$ are the same for all $a,b\in A$. The latter is vacuous when $A$ is almost trivial.
By Proposition 2.7, the vector space $kA$ is a $kA$-module, where $(A,\cdot)$ and $(A,\circ)$ act on $A$, respectively, via left multiplications with respect to $\cdot$ and $\circ$. We shall refer to it as the left regular module over $kA$.
The following are counterparts of Definitions 2.9 and 2.10.
Definition 5.4.
Let $V_{1}$ and $V_{2}$ be $kA$-modules.
  1. (1)
    Their direct sum is defined to be the vector space $V_{1}\oplus V_{2}$, where
    $a\cdot(v_{1},v_{2})$ $=(a\cdot v_{1},a\cdot v_{2})$
    $a\circ(v_{1},v_{2})$ $=(a\circ v_{1},a\circ v_{2})$
    for all $a\in A$ and $v_{1}\in V_{1},\,v_{2}\in V_{2}$.
  2. (2)
    Their tensor product is defined to be the vector space $V_{1}\otimes_{k}V_{2}$, where
    $a\cdot(v_{1}\otimes v_{2})$ $=(a\cdot v_{1})\otimes(a\cdot v_{2})$
    $a\circ(v_{1}\otimes v_{2})$ $=(a\circ v_{1})\otimes(a\circ v_{2})$
    for all $a\in A$ and $v_{1}\in V_{1},\,v_{2}\in V_{2}$.
For both constructions, the relation (5.1) is clearly satisfied.
Definition 5.5.
Let $V_{1}$ and $V_{2}$ be $kA$-modules. A homomorphism from $V_{1}$ to $V_{2}$ is a linear map $f:V_{1}\longrightarrow V_{2}$ such that both of the conditions
$f(a\cdot v)$ $=a\cdot f(v)$
$f(a\circ v)$ $=a\circ f(v)$
hold for all $a\in A$ and $v\in V$. An isomorphism is a bijective homomorphism. We shall say that $V_{1}$ and $V_{2}$ are isomorphic, denoted by $V_{1}\simeq V_{2}$, if there is an isomorphism between them.
Under the correspondence (5.2), clearly two $kA$-modules are isomorphic if and only if the associated representations of $A$ are equivalent.
Definition 5.6.
Let $V$ be a $kA$-module. A subset $U$ of $V$ is a submodule if it is both a $k(A,\cdot)$-submodule and a $k(A,\circ)$-submodule. In this case, clearly $U$ is also a $kA$-module under the same actions. We shall say that $V$ is
  1. (1)
    simple if $V\neq\{0\}$ and $V$ does not have any non-zero proper submodule;
  2. (2)
    indecomposable if $V\neq\{0\}$ and $V$ cannot be expressed as a direct sum of two non-zero proper submodules;
  3. (3)
    semisimple if $V$ can be expressed as a direct sum of simple submodules.
Under the correspondence (5.2), clearly “simple”, “indecomposable”, and “semisimple” for $kA$-modules, respectively, correspond to “irreducible”, “indecomposable”, and “completely reducible” for representations of $A$.
The following are restatements of Proposition 3.2 and Corollary 3.3.
Proposition 5.7.
Let $V$ be a $kA$-module.
  1. $(a)$
    If $U$ is a $k(A,\cdot)$-submodule of $V$, then so is $a\circ U$ for any $a\in A$, and in particular, we get that the sum
    $\sum_{a\in A}a\circ U$
    is a $kA$-submodule of $V$.
  2. $(b)$
    If $U$ is a $k(A,\circ)$-submodule of $V$, then the sum
    $\sum_{b\in A}b\cdot U$
    is a $kA$-submodule of $V$.
Corollary 5.8.
Let $V$ be a $kA$-module. The sums
$\sum_{a,b\in A}a\circ(b\cdot U)\mbox{ and }\sum_{a,b\in A}b\cdot(a\circ U)$
are $kA$-submodules of $V$ for any subspace $U$ of $V$.
An advantage of employing the module-theoretic point of view is that we can further apply Proposition 5.7 to prove the following. It makes use of the well-known fact that over a ring, the sum of finitely many simple submodules is semisimple (see [12, Lemma 1.2.3]).
Corollary 5.9.
Let $A=(A,\cdot,\circ)$ be a finite skew brace. Then every simple $kA$-module is semisimple as a $k(A,\cdot)$-module.
Proof.
Let $S$ be a simple $kA$-module and let $U$ be a simple $k(A,\cdot)$-submodule of $S$. Such a $U$ must exist because $\dim_{k}(S)$ is finite by Proposition 3.4. From Proposition 5.7(a) and the simplicity of $S$, we deduce that
$S=\sum_{a\in A}a\circ U,$
where each summand $a\circ U$ is a $k(A,\cdot)$-submodule. For any $a\in A$, if $W$ is a $k(A,\cdot)$-submodule of $a\circ U$, then $\overline{a}\circ W$ is a $k(A,\cdot)$-submodule of $U$ again by Proposition 5.7(a), so we see that $a\circ U$ is also a simple $k(A,\cdot)$-module. This implies that $S$ is a sum of finitely many simple $k(A,\cdot)$-modules, whence $S$ is semisimple as a $k(A,\cdot)$-module. ∎
Remark 5.10.
For a finite skew brace $A=(A,\cdot,\circ)$, a simple $kA$-module need not be semisimple as a $k(A,\circ)$-module in general (c.f. Example 3.7).
The analog of the aforementioned [12, Lemma 1.2.3] holds for $kA$-modules essentially by the exact same proof. We simply have to note the obvious fact that $kA$-submodules are closed under sums and intersections.
Proposition 5.11.
Let $V=U_{1}+\cdots+U_{n}$ be a $kA$-module that is the sum of finitely many simple $kA$-submodules $U_{1},\dots,U_{n}$. Then
$V=\bigoplus_{i\in I}U_{i}$
for some subset $I$ of $\{1,\dots,n\}$, so in particular $V$ is semisimple.
Proof.
Take $I$ to be an element of maximal size of the set
$\Bigg{\{}J:J\mbox{ is a subset of $\{1,\dots,n\}$ such that the sum $\sum\limits_{j\in J}U_{j}$ is direct}\Bigg{\}},$
which is non-empty because it contains the singletons $\{i\}$. Put $U=\bigoplus_{i\in I}U_{i}$. Then $U=V$, for otherwise there would exist a $j\notin I$ such that $U_{j}\not\subseteq U$, and so $U_{j}\cap U=\{0\}$ because $U_{j}\cap U$ is a proper submodule of the simple module $U_{j}$. But this would imply that $\sum_{i\in I\cup\{j\}}U_{i}$ is a direct sum, which contradicts the maximality of $I$. This completes the proof. ∎
Finally, we prove an analog of the Krull-Schmidt theorem for $kA$-modules. Again, the proof is very similar to that for modules over a ring, but we shall first need a couple lemmas. For a $kA$-module $V$, let
$\mathrm{End}_{kA}(V)=\mathrm{End}_{k(A,\cdot)}(V)\cap\mathrm{End}_{k(A,\circ)}(V)$
denote the set consisting of all $kA$-endomorphisms of $V$.
Analogous to group representations, we have the following lemmas.
Lemma 5.12.
For any $kA$-module $V$, the following are equivalent:
  1. $(1)$
    $V$ is an indecomposable $kA$-module;
  2. $(2)$
    $\mathrm{End}_{kA}(V)$ has no idempotent other than $0_{V}$ and $\mathrm{id}_{V}$.
Proof of $(1)\Rightarrow(2)$.
Suppose that $\pi$ is an idempotent of $\mathrm{End}_{kA}(V)$. Then
$V=\ker(\pi)\oplus\mathrm{Im}(\pi),$
where $\ker(f)$ and $\mathrm{Im}(f)$ are $kA$-submodules of $V$. Thus, it follows from (1) that either $\pi=0_{V}$ or $\pi$ is an isomorphism. In the latter case, note that we must have $\pi=\mathrm{id}_{V}$ because $\pi^{2}=\pi$. ∎
Proof of $(2)\Rightarrow(1)$.
Suppose that $V=U\oplus U^{\prime}$ for $kA$-submodules $U$ and $U^{\prime}$ of $V$. Then projection onto $U$ yields an idempotent $\pi$ of $\mathrm{End}_{kA}(V)$, and (2) implies that either $U=\{0\}$ or $U=V$, whence $V$ is indecomposable. ∎
Lemma 5.13.
For any indecomposable $kA$-module $V$ such that $\dim_{k}(V)$ is finite, the endomorphism ring $\mathrm{End}_{kA}(V)$ is local.
Proof.
Plainly $\mathrm{End}_{kA}(V)$ is a $k$-subalgebra of $\mathrm{End}_{k}(V)$, and is finite dimensional as a $k$-vector space since $V$ is finite dimensional, so we can apply [12, Proposition 11.1.4(2)]. Since $\mathrm{End}_{kA}(V)$ has no idempotent other than $0_{V}$ and $\mathrm{id}_{V}$ by Lemma 5.12, we deduce that $\mathrm{End}_{kA}(V)$ is a local ring. ∎
We are now ready to prove the analog of Krull-Schmidt Theorem for $kA$-modules. Lemma 5.13 is important for the proof because we shall need the standard fact that in a local ring, the sum of any non-invertible elements is again non-invertible (see [12, Proposition 11.1.3] for example).
Theorem 5.14 (Analog of Krull-Schmidt theorem).
Let $V$ be a $kA$-module such that $\dim_{k}(V)$ is finite. Let
$V$ $=U_{1}\oplus\cdots\oplus U_{r}$
$V$ $=U_{1}^{\prime}\oplus\cdots\oplus U_{s}^{\prime}$
be two decompositions of $V$, where $U_{1},\dots,U_{r},U_{1}^{\prime},\dots,U_{s}^{\prime}$ are indecomposable $kA$-submodules of $V$. Then $s=r$ holds, and the $U_{1}^{\prime},\dots,U_{r}^{\prime}$ may be reordered in a way such that $U_{i}\simeq U_{i}^{\prime}$ for all $1\leq i\leq r$.
Proof.
For each $1\leq i\leq r$ and $1\leq j\leq s$, let
$\begin{cases}\iota_{i}:U_{i}\longrightarrow V\\ \iota_{j}^{\prime}:U_{j}^{\prime}\longrightarrow V\end{cases}\mbox{and }\,\ \begin{cases}\pi_{i}:V\longrightarrow U_{i}\\ \pi_{j}^{\prime}:V\longrightarrow U_{j}^{\prime}\end{cases}$
denote the natural inclusion and projection maps, respectively. They are all $kA$-homomorphisms because the $U_{\ell}$ and $U_{\ell}^{\prime}$ are $kA$-submodules. Note that
$\mathrm{id}_{U_{1}}=\pi_{1}\circ\mathrm{id}_{V}\circ\iota_{1}=\sum_{j=1}^{s}(\pi_{1}\circ\iota_{j}^{\prime}\circ\pi_{j}^{\prime}\circ\iota_{1}).$
Since $\mathrm{End}_{kA}(U_{1})$ is a local ring by Lemma 5.13, one of the summands must be invertible. We may assume that $\pi_{1}\circ\iota_{1}^{\prime}\circ\pi_{1}^{\prime}\circ\iota_{1}$ is invertible. Put
$f=\pi_{1}^{\prime}\circ\iota_{1},\,\ g=\pi_{1}\circ\iota_{1}^{\prime},\,\ \theta=g\circ f.$
which are $kA$-homomorphisms. Since $\theta$ is invertible, the map $g$ is surjective and we have an exact sequence
[…]
of $k$-modules that is right split by $f\circ\theta^{-1}$, so then
$U_{1}^{\prime}=\ker(g)\oplus\mathrm{Im}(f\circ\theta^{-1})=\ker(g)\oplus\mathrm{Im}(f).$
Since the summands are $kA$-submodules and $U_{1}^{\prime}$ is indecomposable, we get
$\ker(g)=\{0\}\mbox{ and }\mathrm{Im}(f)=U_{1}^{\prime}.$
Note that the other way round is impossible because $g\circ f$ is invertible. We deduce that $g$ is bijective, whence $U_{1}\simeq U_{1}^{\prime}$. Now, we have
$\bigoplus_{\ell=2}^{r}U_{\ell}\simeq\bigoplus_{\ell=2}^{s}U_{\ell}^{\prime}\mbox{ via }u\mapsto\sum_{i=2}^{r}\pi_{\ell}^{\prime}(u),$
where injectivity holds because the injectivity of $g$ implies that
$U_{1}^{\prime}\cap\bigoplus_{\ell=2}^{r}U_{\ell}=\{0\},$
and then surjectivity follows by comparing dimensions. Hence, by induction on $\min\{r,s\}$, we deduce that $s=r$, and the $U_{2}^{\prime},\dots,U_{r}^{\prime}$ may be reordered in a way such that $U_{i}\simeq U_{i}^{\prime}$ for all $2\leq i\leq r$. Since we have already shown that $U_{1}\simeq U_{1}^{\prime}$, the claim follows. ∎
6. Analog of Clifford’s theorem
Let us first recall Clifford’s theorem for group representations.
Theorem 6.1 (Clifford’s theorem).
Let $G$ be a finite group and let $N$ be a normal subgroup of $G$. The restriction $\mathrm{Res}_{N}^{G}(S)$ of any simple $kG$-module $S$ to $kN$ is always semisimple as a $kN$-module.
Proof.
See [12, Theorem 5.3.1]. ∎
The analog of Clifford’s theorem for skew brace representations, where we use “ideal” as the analog of “normal subgroup”, is true. To prove it, we shall need the following lemma.
Lemma 6.2.
Let $A=(A,\cdot,\circ)$ be any skew brace and let $V$ be a $kA$-module. Let $I$ be an ideal of $A$ and let $U$ be a $kI$-submodule of $V$. For any $a\in A$, we have that $a\cdot U$ and $a\circ U$ are also $kI$-submodules of $V$. Moreover, if $U$ is a simple $kI$-module, then so are $a\cdot U$ and $a\circ U$.
Proof.
Let $u\in U$ and $x\in I$. Since $I$ is normal in both $(A,\cdot)$ and $(A,\circ)$, we have $a^{-1}\cdot x\cdot a,\,\overline{a}\circ x\circ a\in I$. Then
$x\cdot(a\cdot u)$ $=a\cdot((a^{-1}\cdot x\cdot a)\cdot u)\in a\cdot U$
$x\circ(a\circ u)$ $=a\circ((\overline{a}\circ x\circ a)\circ u)\in a\circ U$
since $U$ is a $kI$-submodule. On the other hand, by applying the relation (5.1) with $(a,b,v)$ replaced by $(x,a,x\circ u)$, we obtain
$x\circ(a\cdot u)$ $=\lambda_{x}^{\mbox{\tiny op}}(a)\cdot(x\circ u)$
$=((a\circ\overline{a}\circ x\circ a)\cdot x^{-1})\cdot(x\circ u)$
$=a\cdot(\lambda_{a}(\overline{a}\circ x\circ a)\cdot x^{-1}\cdot(x\circ u)).$
Since $I$ is an ideal of $A$, we have $\lambda_{a}(\overline{a}\circ x\circ a)\in I$ and so $x\circ(a\cdot u)\in a\cdot U$. Similarly, by applying (5.1) with $(a,b,v)$ replaced by $(\overline{a},x,u)$, we obtain
$x\cdot(a\circ u)$ $=a\circ(\lambda_{\overline{a}}^{\mbox{\tiny op}}(x)\cdot u)$
$=a\circ((\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1})\cdot u).$
Since $I$ is an ideal of $A$, we have $\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1}\in I$ and so $x\cdot(a\circ u)\in a\circ U$. Thus, indeed $a\cdot U$ and $a\circ U$ are $kI$-submodules of $V$.
For any $kI$-submodule $W$ of $V$, observe that
$W\subseteq a\cdot U\implies a^{-1}\cdot W\subseteq U,\,\ W\subseteq a\circ U\implies\overline{a}\circ W\subseteq U.$
Since $a^{-1}\cdot W$ and $\overline{a}\circ W$ are $kI$-submodules by the first claim, we see that if $U$ is simple as a $kI$-module, then so are $a\cdot U$ and $a\circ U$.∎
Theorem 6.3 (Analog of Clifford’s theorem).
Let $A=(A,\cdot,\circ)$ be a finite skew brace and let $I$ be an ideal of $A$. The restriction $\mathrm{Res}_{I}^{A}(S)$ of any simple $kA$-module $S$ to $kI$ is always semisimple as a $kI$-module.
Proof.
Let $U$ be any simple $kI$-submodule of $\mathrm{Res}_{I}^{A}(S)$. Such a $U$ must exist because $\dim_{k}(S)$ is finite by Proposition 3.4. Since $S$ is a simple $kA$-module, we deduce from Corollary 5.8 that
$S=\sum_{a,b\in A}a\circ(b\cdot U)$
must hold. Each summand $a\circ(b\cdot U)$ is a simple $kI$-submodule by Lemma 6.2. It then follows from Proposition 5.11 that the $kI$-module $\mathrm{Res}_{I}^{A}(S)$, being a sum of finitely many simple $kI$-submodules, is semisimple. ∎
In the setting of Theorem 6.1, we may write
(6.1) $\mathrm{Res}_{N}^{G}(S)=\bigoplus_{i=1}^{r}S_{i}^{\oplus{m_{i}}},$
where $r\in\mathbb{N}$, and $S_{1},\dots,S_{r}$ are pairwise non-isomorphic simple $kN$-modules occurring with multiplicities $m_{1},\dots,m_{r}$. We shall refer to $S_{1}^{\oplus m_{1}},\dots,S_{r}^{\oplus m_{r}}$ as the homogeneous components of $\mathrm{Res}_{N}^{G}(S)$. There is in fact a stronger form of Clifford’s theorem that gives information about the decomposition (6.1).
Theorem 6.4 (Clifford’s theorem, strong form).
Let $G$ be a finite group and let $N$ be a normal subgroup of $G$. For any simple $kG$-module $S$, we have:
  1. $(a)$
    $G$ permutes the homogeneous components transitively;
  2. $(b)$
    $m_{1}=\cdots=m_{r}$ and $\dim_{k}(S_{1})=\cdots=\dim_{k}(S_{r})$;
in the decomposition (6.1).
Proof.
See [12, Theorem 5.3.2]. ∎
In the setting of Theorem 6.3, analogous to (6.1) we may write
(6.2) $\mathrm{Res}_{I}^{A}(S)=\bigoplus_{i=1}^{r}S_{i}^{\oplus{m_{i}}},$
where $r\in\mathbb{N}$, and $S_{1},\dots,S_{r}$ are pairwise non-isomorphic simple $kI$-modules occurring with multiplicities $m_{1},\dots,m_{r}$. We shall refer to $S_{1}^{\oplus m_{1}},\dots,S_{r}^{\oplus m_{r}}$ as the homogeneous components of $\mathrm{Res}_{I}^{A}(S)$. Naturally, we wonder whether the analog of (b) and some sort of analog of (a) in Theorem 6.4 are true.
Lemma 6.5.
Let $A=(A,\cdot,\circ)$ be any skew brace and let $V$ be a $kA$-module. Let $I$ be an ideal of $A$ and let $U_{1},\,U_{2}$ be $kI$-submodules of $V$. For any $a\in A$, as $kI$-modules, we have
  1. $(a)$
    $U_{1}\simeq U_{2}$ implies $a\cdot U_{1}\simeq a\cdot U_{2}$ and $a\circ U_{1}\simeq a\circ U_{2}$;
  2. $(b)$
    $a\cdot U_{1}\simeq a\cdot U_{2}$ or $a\circ U_{1}\simeq a\circ U_{2}$ implies $U_{1}\simeq U_{2}$.
Proof of $(a)$.
Let $f:U_{1}\longrightarrow U_{2}$ be any $kI$-module isomorphism. Consider
$\begin{cases}f_{\bullet}:a\cdot U_{1}\longrightarrow a\cdot U_{2};&f_{\bullet}(a\cdot u)=a\cdot f(u)\,\ (u\in U_{1}),\\ f_{\circ}:a\circ U_{1}\longrightarrow a\circ U_{2};&f_{\circ}(a\circ u)=a\circ f(u)\,\ (u\in U_{1}),\end{cases}$
which are easily seen to be $k$-isomorphisms. Let $u\in U_{1}$ and $x\in I$. Since $I$ is normal in both $(A,\cdot)$ and $(A,\circ)$, we have $a^{-1}\cdot x\cdot a,\,\overline{a}\circ x\circ a\in I$. Then
$f_{\bullet}(x\cdot(a\cdot u))$ $=a\cdot f((a^{-1}\cdot x\cdot a)\cdot u)$
$=a\cdot(a^{-1}\cdot x\cdot a)\cdot f(u)$
$=x\cdot f_{\bullet}(a\cdot u),$
$f_{\circ}(x\circ(a\circ u))$ $=a\circ f((\overline{a}\circ x\circ a)\circ u)$
$=a\circ(\overline{a}\circ x\circ a)\circ f(u)$
$=x\circ f_{\circ}(a\circ u),$
because $f$ is a $kI$-module homomorphism.
By the proof of Lemma 6.2, for any $v\in U_{1}\cup U_{2}$, we have
(6.3) $\begin{cases}x\circ(a\cdot v)=a\cdot(\lambda_{a}(\overline{a}\circ x\circ a)\cdot x^{-1}\cdot(x\circ v)),\\ x\cdot(a\circ v)=a\circ((\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1})\cdot v),\end{cases}$
and $\lambda_{a}(\overline{a}\circ x\circ a),\,\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1}\in I$ because $I$ is an ideal of $A$. From these two identities and the fact that $f$ is a $kI$-homomorphism, we see that
$f_{\bullet}(x\circ(a\cdot u))$ $=a\cdot f(\lambda_{a}(\overline{a}\circ x\circ a)\cdot x^{-1}\cdot(x\circ u))$
$=a\cdot(\lambda_{a}(\overline{a}\circ x\circ a)\cdot x^{-1}\cdot(x\circ f(u)))$
$=x\circ(a\cdot f(u))$
$=x\circ f_{\bullet}(a\cdot u),$
where we have applied (6.3) in the first and third equalities. Similarly
$f_{\circ}(x\cdot(a\circ u))$ $=a\circ f((\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1})\cdot u)$
$=a\circ((\overline{a}\cdot\lambda_{\overline{a}}(x)\cdot\overline{a}^{-1})\cdot f(u))$
$=x\cdot(a\circ f(u))$
$=x\cdot f_{\circ}(a\circ u).$
Thus, both $f_{\bullet}$ and $f_{\circ}$ are $kI$-isomorphisms. This proves that $a\cdot U_{1}\simeq a\cdot U_{2}$ and $a\circ U_{1}\simeq a\circ U_{2}$ as $kI$-modules, as desired. ∎
Proof of $(b)$.
For any $kI$-submodule $U$ of $V$, we have
$U=a^{-1}\cdot(a\cdot U)\mbox{ and }U=\overline{a}\circ(a\circ U),$
so the claim follows immediately from (a). ∎
Theorem 6.6 (Analog of Clifford’s theorem, strong form).
Let $A=(A,\cdot,\circ)$ be a finite skew brace and let $I$ be an ideal of $A$. For any simple $kA$-module $S$, we have:
  1. $(a)$
    $(A,\cdot)$ and $(A,\circ)$ permute the homogeneous components, and together they permute the homogeneous components transitively, in the sense that
    1. $\bullet$
      for any $a\in A$ and $1\leq i\leq r$, there exist $1\leq j_{\bullet},\,j_{\circ}\leq r$ such that
      $a\cdot S_{i}^{\oplus m_{i}}\simeq S_{j_{\bullet}}^{\oplus m_{j_{\bullet}}}\mbox{ and }a\circ S_{i}^{\oplus m_{i}}\simeq S_{j_{\circ}}^{\oplus m_{j_{\circ}}};$
    2. $\bullet$
      for any $1\leq i,\,j\leq r$, there exist $a,b,c,d\in A$ such that
      $a\circ(b\cdot S_{i}^{\oplus m_{i}})\simeq S_{j}^{\oplus m_{j}}\mbox{ and }c\cdot(d\circ S_{j}^{\oplus m_{j}})\simeq S_{i}^{\oplus m_{i}};$
  2. $(b)$
    $m_{1}=\cdots=m_{r}$ and $\dim_{k}(S_{1})=\cdots=\dim_{k}(S_{r})$;
in the decomposition (6.2).
Proof.
Let $a\in A$. Observe that (6.2) implies
$\mathrm{Res}_{I}^{A}(S)$ $=\bigoplus_{\ell=1}^{r}a\cdot S_{\ell}^{\oplus m_{\ell}}=\bigoplus_{\ell=1}^{r}(a\cdot S_{\ell})^{\oplus m_{\ell}},$
$\mathrm{Res}_{I}^{A}(S)$ $=\bigoplus_{\ell=1}^{r}a\circ S_{\ell}^{\oplus m_{\ell}}=\bigoplus_{\ell=1}^{r}(a\circ S_{\ell})^{\oplus m_{\ell}},$
where $a\cdot S_{\ell}$ and $a\circ S_{\ell}$ are simple $kI$-modules by Lemma 6.2, and
$a\cdot S_{\ell}\not\simeq a\cdot S_{\ell^{\prime}}\mbox{ and }a\circ S_{\ell}\not\simeq a\circ S_{\ell^{\prime}}\mbox{ for all }1\leq\ell<\ell^{\prime}\leq r$
by Lemma 6.5. Notice that $\dim_{k}(S)$ is finite by Proposition 3.4 so Theorem 5.14 applies. By comparing the above decompositions with (6.2), we deduce that the first claim in (a) holds.
Let $1\leq i\leq r$ be fixed. Since $S$ is a simple $kA$-module, it is a consequence of Corollary 5.8 that
$S=\sum_{a,b\in A}a\circ(b\cdot S_{i}^{\oplus m_{i}})=\sum_{a,b\in A}(a\circ(b\cdot S_{i}))^{\oplus m_{i}}.$
Each summand $a\circ(b\cdot S_{i})$ is a simple $kI$-module by Lemma 6.2, so we may refine the above sum to a direct sum by Proposition 5.11. For any $1\leq j\leq r$, again by applying Theorem 5.14, we deduce from (6.2) that
(6.4) $a\circ(b\cdot S_{i})\simeq S_{j},\mbox{ so in particular }\dim_{k}(S_{i})=\dim_{k}(S_{j}),$
for some $a,b\in A$. On the other hand, as above we have
$\mathrm{Res}_{I}^{A}(S)=\bigoplus_{\ell=1}^{r}a\circ(b\cdot S_{\ell}^{\oplus m_{\ell}})=\bigoplus_{\ell=1}^{r}(a\circ(b\cdot S_{\ell}))^{\oplus m_{\ell}},$
where each $a\circ(b\cdot S_{\ell})$ is a simple $kI$-module by Lemma 6.2, and
$a\circ(b\cdot S_{\ell})\not\simeq a\circ(b\cdot S_{\ell^{\prime}})\mbox{ for all $1\leq\ell<\ell^{\prime}\leq r$}$
by Lemma 6.5. Thus, by comparing the above decomposition with (6.2), we see from (6.4) and Theorem 5.14 that $m_{i}=m_{j}$. It follows that
$a\circ(b\cdot S_{i}^{\oplus m_{i}})\simeq(a\circ(b\cdot S_{i}))^{\oplus m_{i}}\simeq S_{j}^{\oplus m_{j}}.$
Note that then Lemma 6.5 implies that
$S_{i}^{\oplus m_{i}}\simeq b^{-1}\cdot(\overline{a}\circ S_{j}^{\oplus m_{j}}),$
so we simply take $c=b^{-1}$ and $d=\overline{a}$. This proves the second claim in (a).
In the above, we have shown that $\dim_{k}(S_{i})=\dim_{k}(S_{j})$ and $m_{i}=m_{j}$ for any arbitrary $1\leq i,\,j\leq r$, so the claim (b) holds as well. ∎
Let us conclude by remarking that $(A,\cdot)$ on its own does not permute the homogeneous components transitively in general, and similarly for $(A,\circ)$. To give an explicit example, consider an almost trivial skew brace $A=(A,\cdot,\circ)$, so $\circ$ is the opposite operation $\cdot^{\mbox{\tiny op}}$ of $\cdot$. Note that $(A,\cdot)\simeq(A,\circ)$, and ideals of $A$ coincide with normal subgroups of $(A,\cdot)$ or equivalently normal subgroups of $(A,\circ)$. Let {$\{\star,\star^{\prime}\}=\{\cdot,\circ\}$ and take $S$ to be any simple $k(A,\star)$-module.
  1. (1)
    Regard $S$ as a $kA$-module by letting $(A,\star^{\prime})$ act trivially (recall Examples 5.2 and 5.3). Since the action of $(A,\star^{\prime})$ is trivial, the notions of $k(A,\star)$-submodules and $kA$-submodules coincide here.
  2. (2)
    Suppose that $I$ is an ideal of $A$ and
    $\mathrm{Res}_{I}^{A}(S)=\mathrm{Res}_{(I,\star)}^{(A,\star)}(S)=\bigoplus_{i=1}^{r}S_{i}^{\oplus{m_{i}}},$
    where $S_{1},\dots,S_{r}$ are pairwise non-isomorphic simple $k(I,\star)$-modules. By letting $(I,\star^{\prime})$ act trivially, they are also simple $kI$-modules. In the case that $r\geq 2$, the group $(A,\star^{\prime})$ cannot permute the homogeneous components transitively because its action is trivial.
In conclusion, it suffices to give an example of a finite group $G$ and a simple $kG$-module $S$ for which $\mathrm{Res}_{N}^{G}(S)$ has at least two homogeneous components for some normal subgroup $N$ of $G$.
Example 6.7.
Assume $\mathrm{char}(k)=2$ and let $\omega\in k$ be a primitive cube root of unity (assume it exists). Let $G=S_{4}$ and $N=A_{4}$. For the unique simple $kS_{4}$-module $S$ of degree $2$, the $kA_{4}$-module
$\mathrm{Res}_{A_{4}}^{S_{4}}(S)=U_{1}\oplus U_{2}$
has two homogeneous components. Here the actions of $A_{4}$ on $U_{1}$ and $U_{2}$ both factor through the Klein four-subgroup of $A_{4}$. Moreover, the element (1 2 3)123(1\,2\,3)( 1 2 3 ) acts as $\omega$ on $U_{1}$ and as $\omega^{2}$ on $U_{2}$, so that $U_{1}\not\simeq U_{2}$ as $kA_{4}$-modules.
7. Simple modules in prime characteristic
For a group $G$, the trivial module over $kG$, denoted by $k_{G}$, is defined to be the one-dimensional vector space over $k$ on which $G$ acts trivially. We have the following well-known result from modular representation theory.
Theorem 7.1.
Let $G$ be a finite group and let $k$ be a field of prime characteristic $p$. The following are equivalent.
  1. $(1)$
    $|G|$ is a power of $p$;
  2. $(2)$
    $k_{G}$ is the only simple $kG$-module.
Proof of $(1)\Rightarrow(2)$.
We shall give a proof since we need it for later. Let $S$ be a simple $kG$-module and let $u\in S$ be any non-zero vector. The $\mathbb{F}_{p}$-subspace $U$ generated by $\{g\cdot u:g\in G\}$ is clearly an $\mathbb{F}_{p}G$-submodule and consider the action of $G$ on $U$. An orbit must have size $1$ or a power of $p$ because $G$ is a $p$-group. There is at least one fixed point, namely the zero vector. Since $|U|$ is a power of $p$, it follows from the class equation that there must be another fixed point $v\in U\setminus\{0\}$. Then $\mathrm{span}_{k}(v)$ is a non-zero $kG$-submodule on which $G$ acts trivially. Since $S$ is simple, we have $S=\mathrm{span}_{k}(v)$, and so $S$ is a trivial $kG$-module.∎
Proof of $(2)\Rightarrow(1)$.
We give a proof for the convenience of the reader. Let $P$ be any Sylow $p$-subgroup of $G$. The induced module $kG\otimes_{kP}k_{P}$ of $k_{P}$ has $k_{G}$ as a direct summand (see [2, p. 67–68]), say
$kG\otimes_{kP}k_{P}=k_{G}\oplus X,$
where $X$ is a $kG$-module. Note that
$\mathrm{Hom}_{kG}(k_{G},kG\otimes_{kP}k_{P})\simeq\mathrm{Hom}_{kP}(k_{P},k_{P})$
by Frobenius reciprocity (see [12, Corollary 4.3.8]). Since
$\mathrm{Hom}_{kG}(k_{G},k_{G}\oplus X)\simeq\mathrm{Hom}_{kG}(k_{G},k_{G})\oplus\mathrm{Hom}_{kG}(k_{G},X),$
by comparing dimensions, we see that
$\dim_{k}(\mathrm{Hom}_{kG}(k_{G},X))=0,$
which implies that $X$ does not have $k_{G}$ as a $kG$-submodule. But this is only possible when $X=\{0\}$ because any simple $kG$-submodule of $X$ is trivial by the hypothesis. Then $kG\otimes_{kP}k_{P}$ is the trivial $kG$-module $k_{G}$, and since
$\dim_{k}(kG\otimes_{kP}k_{P})=[G:P]\dim_{k}(k_{P})=[G:P],$
we get that $G=P$ and so $G$ is a $p$-group. ∎
For a skew brace $A=(A,\cdot,\circ)$, we similarly define the trivial module over $kA$, denoted by $k_{A}$, to be the one-dimensional vector space over $k$ on which both $(A,\cdot)$ and $(A,\circ)$ act trivially. The analog of Theorem 7.1 holds for skew brace representations, as follows.
Theorem 7.2.
Let $A=(A,\cdot,\circ)$ be a finite skew brace and let $k$ be a field of prime characteristic $p$. The following are equivalent.
  1. $(1)$
    $|A|$ is a power of $p$;
  2. $(2)$
    $k_{A}$ is the only simple $kA$-module.
Proof of $(1)\Rightarrow(2)$.
Let $S$ be a simple $kA$-module. Since $\dim_{k}(S)$ is finite by Proposition 3.4, it must contain a simple $k(A,\cdot)$-submodule $S_{\bullet}$. Let $u\in S_{\bullet}$ be any non-zero vector. By considering the action of $(A,\circ)$ on the $\mathbb{F}_{p}$-subspace $U$ generated by $\{a\circ u:a\in A\}$, we deduce as in the Theorem 7.1 that there is a non-zero $v\in U$ that is fixed by the action of $(A,\circ)$.
Now, note that $S_{\bullet}$ must be the trivial $k(A,\cdot)$-module by Theorem 7.1. For each $a\in A$, since $u\in S_{\bullet}$, the relation (5.1) then implies that
$\lambda_{a}^{\mbox{\tiny op}}(b)\cdot(a\circ u)$ $=a\circ(b\cdot(\overline{a}\circ(a\circ u)))$
$=a\circ(b\cdot u)$
$=a\circ u$
for all $b\in A$. Since $\lambda_{a}^{\mbox{\tiny op}}$ is a bijection on $A$, this shows that $a\circ u$ is fixed by the action of $(A,\cdot)$. Since $v$ is a linear combination of such elements $a\circ u$, it follows that $v$ is also fixed by the action of $(A,\cdot)$.
We have thus found a non-zero $v$ that is fixed by the actions of both $(A,\cdot)$ and $(A,\circ)$. Then $\mathrm{span}_{k}(v)$ is a non-zero $kA$-submodule on which both $(A,\cdot)$ and $(A,\circ)$ act trivially. Since $S$ is simple, we have $S=\mathrm{span}_{k}(v)$, and so $S$ is a trivial $kA$-module. ∎
Proof of $(2)\Rightarrow(1)$.
Consider the contrapositive and assume that $|A|$ is not a power of $p$. By Theorem 7.1, there exists a non-trivial simple $k(A,\circ)$-module $S$, and regard $S$ as a $kA$-module by letting $(A,\cdot)$ act trivially (recall Example 5.2). Then, clearly $S$ is non-trivial and simple as a $kA$-module as well. ∎
In the setting of Theorem 7.2, where we assume that $|A|$ is not a power of $p$ (the characteristic of the field $k$), in certain cases it is possible to exhibit simple $kA$-modules for which both the actions of $(A,\cdot)$ and $(A,\circ)$ are non-trivial. To that end, consider the group
$A/A*A=(A/A*A,\cdot)=(A/A*A,\circ),$
and let $S$ be any non-trivial one-dimensional $k(A/A*A)$-module (assume it exists). For example, such an $S$ always exists when $A/A*A$ is an abelian group whose order is not a power of $p$ and $k$ is a splitting field for $A/A*A$.
  1. (1)
    Let $S_{\circ}=S$ be the $kA$-module on which $(A,\circ)$ acts via the quotient map $(A,\circ)\longrightarrow(A/A*A,\circ)$ and the given $A/(A*A)$-action, while $(A,\cdot)$ acts trivially (recall Example 5.2).
  2. (2)
    Let $S_{\bullet}=S$ be the $kA$-module on which $(A,\cdot)$ acts via the quotient map $(A,\cdot)\longrightarrow(A/A*A,\cdot)$ and the given $A/(A*A)$-action, while $(A,\circ)$ acts trivially (recall Example 5.3; for any $a,b\in A$, note that
    $\lambda_{a}^{\mbox{\tiny op}}(b)$ $\equiv a\cdot b\cdot a^{-1}\pmod{A*A}$
    $\equiv b\pmod{(A*A)[A,A]},$
    where $[A,A]$ denotes the commutator subgroup of $(A,\cdot)$. The action of $(A,\cdot)$ must also factor through $[A,A]$ because $S_{\bullet}$ is one-dimensional, and
    this means that $\lambda_{a}^{\mbox{\tiny op}}(b)$ and $b$ have the same actions.)
Consider their tensor product $S_{\bullet}\otimes_{k}S_{\circ}$ defined as in Definition 5.4. Clearly $S_{\bullet}\otimes_{k}S_{\circ}$ is a simple $kA$-module because it is one-dimensional. Moreover, for any $a\in A$, since $S$ is one-dimensional, there exists $\ell_{a}\in k^{\times}$ such that
$a\cdot(s_{\bullet}\otimes s_{\circ})$ $=(a\cdot s_{\bullet})\otimes(a\cdot s_{\circ})$
$=(\ell_{a}s_{\bullet})\otimes s_{\circ}$
$=\ell_{a}(s_{\bullet}\otimes s_{\circ})$
$a\circ(s_{\bullet}\otimes s_{\circ})$ $=(a\circ s_{\bullet})\otimes(a\circ s_{\circ})$
$=s_{\bullet}\otimes(\ell_{a}s_{\circ})$
$=\ell_{a}(s_{\bullet}\otimes s_{\circ})$
for all $s_{\bullet}\in S_{\bullet}$ and $s_{\circ}\in S_{\circ}$. By the non-triviality of $S$, we have $\ell_{a}\neq 1$ for some $a\in A$. This means that the actions of $(A,\cdot)$ and $(A,\circ)$ on the simple $kA$-module $S_{\bullet}\otimes_{k}S_{\circ}$ are both non-trivial.
Example 7.3.
Consider the brace $A=(\mathbb{Z}/r\mathbb{Z}\times\mathbb{Z}/q\mathbb{Z},+,\circ)$, where
$\begin{pmatrix}a_{1}\\ a_{2}\end{pmatrix}\circ\begin{pmatrix}b_{1}\\ b_{2}\end{pmatrix}=\begin{pmatrix}a_{1}+\lambda^{a_{2}}b_{1}\\ a_{2}+b_{2}\end{pmatrix}$
for all $\left(\begin{smallmatrix}a_{1}\\ a_{2}\end{smallmatrix}\right),\left(\begin{smallmatrix}b_{1}\\ b_{2}\end{smallmatrix}\right)\in\mathbb{Z}/r\mathbb{Z}\times\mathbb{Z}/q\mathbb{Z}$ (c.f. [1, Theorem]). Here $r,q$ are primes for which $r\equiv 1\pmod{q}$, and $\lambda\in\mathbb{Z}/r\mathbb{Z}$ is an element of multiplicative order $q$. Observe that $A*A=\mathbb{Z}/r\mathbb{Z}\times\{0\}$ and $A/A*A\simeq\mathbb{Z}/q\mathbb{Z}$. Thus, when $k$ is a splitting field of $A/A*A$ of characteristic coprime to $q$, for example, there is a non-trivial one-dimensional $k(A/A*A)$-module, and we can use the above method to construct a simple $kA$-module for which the actions of both $(A,\cdot)$ and $(A,\circ)$ are non-trivial.
Acknowledgments
This project started from discussions at the symposium (open) “Research on finite groups, algebraic combinatorics, and vertex algebras” held in December 2023 at the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
This work is supported by JSPS KAKENHI Grant Number 24K16891.
The authors would like to thank Kyoichi Suzuki for useful discussions on modular representation theory of finite groups.